A firm is comparing two designs for a new product, designated “Design 1” and “Design 2.” Data concerning the design for disassembly for both designs are provided in the tables below.

Design 1:

Component

Resale Revenue

per Unit

Recycling Revenue

per Unit

Processing Cost

per Unit

Disposal Cost

per Unit

A

$5.60

$8.90

$3.50

$2.85

B

$2.56

$3.67

$3.25

$4.00

C

$1.34

$4.32

$0.69

$3.34

Design 2:

Component

Resale Revenue

per Unit

Recycling Revenue

per Unit

Processing Cost

per Unit

Disposal Cost

per Unit

A

$4.76

$5.68

$6.00

$2.13

B

$3.21

$5.28

$2.42

$3.59

C

$5.31

$6.98

$2.47

$4.90

a)Which design appears to be the best environmental design based on its revenue retrieval?

b)what is the dollar value of the revenue retrieval per unit?
Answer #1

Design 1:

Component

Resale Revenue

per Unit

Recycling Revenue

per Unit

Processing Cost

per Unit

Disposal Cost

per Unit

A

$5.60

$8.90

$3.50

$2.85

B

$2.56

$3.67

$3.25

$4.00

C

$1.34

$4.32

$0.69

$3.34

Total

$9.5

$16.89

$7.44

$10.19

Design 2:

Component

Resale Revenue

per Unit

Recycling Revenue

per Unit

Processing Cost

per Unit

Disposal Cost

per Unit

A

$4.76

$5.68

$6.00

$2.13

B

$3.21

$5.28

$2.42

$3.59

C

$5.31

$6.98

$2.47

$4.90

Total

$13.28

$17.94

$$10.89

$10.62

Revenue retrieval = (Total resale revenue + Total recycling revenue) – (Total processing cost +Total disposal cost)

Design 1

Revenue retrieval = (9.5 + 16.89) – (7.44 + 10.19)

= 26.39 – 17.63

= $8.76

Design 2

Revenue retrieval = (13.28 + 17.94) – (10.89 + 10.62)         

= 31.22 -21.51

= $9.71

a) Design 2 appears to be the best environmental design as revenue retrieval is more of Design 2.

b) The dollar value of the revenue retrieval per unit is $9.71